An analyst has been following American Dream stock. He projects the following dividends for the next three years:

YEAR 1 2 3
Dividend $1.77 $2.18 $2.94

The analyst notes that American Dream stock has a required return of 10.26%. The analyst projects that dividends will grow at a constant rate of 2.00% per year after year 3.

What is the current price of the stock if his assumptions are correct?
Answer #1

P3 = D4 / (k-g)

= 2.94(1.02) / (0.1026 - 0.02)

= 2.9988 / 0.0826

P3 = $36.305

Current Price (P0)

P0 = D1/(1+k) + D2/(1+k)2 + D3/(1+k)3 + P3/(1+k)3

= 1.77/(1.1026) + 2.18/(1.1026)2 + 2.94/(1.1026)3 + 36.305/(1.1026)3

P0 = $32.6757 or $32.68